### v2

Abstract
Over the past decade, visual gaze estimation has garnered growing attention within the research community, thanks to its wide-ranging application scenarios. While existing estimation approaches have achieved remarkable success in enhancing prediction accuracy, they primarily infer gaze directions from single-image signals and discard the huge potentials of the currently dominant text guidance. Notably, visual-language collaboration has been extensively explored across a range of visual tasks, such as image synthesis and manipulation, leveraging the remarkable transferability of large-scale Contrastive Language-Image Pre-training (CLIP) model. Nevertheless, existing gaze estimation approaches ignore the rich semantic cues conveyed by linguistic signals and priors in CLIP feature space, thereby yielding performance setbacks. In pursuit of making up this gap, we delve deeply into the text-eye collaboration protocol and introduce a novel gaze estimation framework in this paper, referred to as GazeCLIP. Specifically, we intricately design a linguistic description generator to produce text signals with coarse directional cues. Additionally, a CLIP-based backbone that excels in characterizing text-eye pairs for gaze estimation is presented. This is followed by the implementation of a fine-grained multi-modal fusion module aimed at modeling the interrelationships between heterogeneous inputs. Extensive experiments on three challenging datasets demonstrate the superiority of the proposed GazeCLIP which surpasses the previous approaches and achieves the state-of-the-art estimation accuracy.
keywords:
[…]
1 Introduction
During the past decade, visual gaze estimation has garnered ever-increasing attention thanks to its widespread application scenarios, such as saliency detection [3, 4], virtual reality[5], human-robot interaction[40, 41], medical diagnosis[42], driver fatigue estimation[6, 7] and other areas[1, 2]. Gaze estimation hammers at inferring accurate gaze direction of a human face or eyes from the input image. While the practical potential of this visual task is undeniably promising, it nonetheless grapples with substantial challenges. For example, the use of diverse camera views inevitably introduces unpredictable variations in lighting conditions, colors, and facial appearances, resulting in severe performance degradation. To struggle with this issue, several existing studies[17][23, 24, 25][51, 52, 53, 54] have made lots of efforts to marry the stability and robustness to regression algorithms. An additional catalyst for the thriving of investigations on gaze estimation is the increasing availability of large-scale datasets [12, 13, 14, 15, 16] which consist of varied subjects including lighting conditions, head poses and so on.
Figure 1: (a) Existing single-modal approaches directly learn gaze-oriented representations from 2D face/eye images via CNNs-based structures, whereas (b) our proposed novel GazeCLIP delves deep into the synergistic effects of text-image features.
Inspired by the huge success of deep learning in computer vision[44, 45, 46, 47], an increasing number of researchers are shifting their focus towards modeling complex mappings between input signals and gaze labels[16, 17][22]. Nonetheless, these methods typically operate on single-modal facial images as input, making them susceptible to the challenging task of extracting gaze-specific features amidst the presence of unrelated regions (e.g., the nose and mouth) scattered throughout the input image as different regions are prone to confuse the model on what to prioritize in learning. A general model for video-based remote POG estimation[8] evades this barrier by mining eye geometric characteristics. Albeit the impressive progress, these conventional geometry-modelling approaches exhibit several drawbacks, such as limited laboratory environment and poor generalization performance to new subject, thereby leading to a surge in the booming of appearance-based gaze estimation algorithms, harnessing the capabilities of cutting-edge deep learning models. Specifically, to enhance the estimation performance, recent learning-based protocols [9, 10, 11] excel in extracting useful features from raw facial or eye images through the design of diverse high-capacity CNN models, see Figure 1(a). Furthermore, the inescapable drastic variations introduced by diverse camera views and individual characteristics (e.g. gender, skin color and glasses) [12][22] easily get the model trapped into failure of learning highly-nonlinear relationships, resulting in overfitting and subpar performance. Therefore, there is a large room for further exploration of gaze regression model with enhanced learning capabilities.
Recently, a cohort of large-scale pre-traiend linguistic-vision models [31, 32, 33, 34] have come to dominate the forefront of artificial intelligence, and trigger the vogue of multi-modal learning. Leveraging copious unlabeled data in conjunction with a contrastive learning framework helps yield robust characteristics. Particularly, Contrastive Language-Image Pre-Training (CLIP) [18] elicits the foremost attention within the research community, and has been adopted by widespread downstream vision tasks, such as image generation and visual question answering [48, 49, 50]. Specifically, CLIP’s backbone is comprised of set of transformer blocks, and is trained using extensive image-text pairs (400 million). A well-trained CLIP empowers the model with fascinating capabilities for modeling inter-relationships between language and visual modalities. The primary advantage of incorporating CLIP lies in its implicit enrichment of features with the nuanced semantics embedded in natural language. For example, object detection [19], semantic segmentation [20], and image classification [21] have achieved notable success by harnessing the guidance ability of the prominent CLIP model.
Despite the remarkable progress attained by previous problems, it is noteworthy that the field of gaze estimation neglects the substantial benefits of leveraging language guidance. Indeed, the task of regressing raw face/eye images into continuous gaze vectors presents a intractable challenge. Therefore, a "divide-and-conquer" idea naturally emerges. It tries to first predict several coarse-grained directions of gaze and then utilize the coarse-granularity pseudo labels to guide the fine-grained gaze estimation (accurate direction vector). Moreover, consider the potential for the gaze estimator to yield less satisfactory results when the target gaze direction deviates from the camera’s central axis, as opposed to cases where the gaze is directly aimed at the camera, we believe that the estimation accuracy can be notably enhanced through the guidance of coarse-grained pseudo directions. Motivated by the booming of CLIP-equipped approaches in other vision tasks, we are pioneering an unprecedented attempt to leverage the collaborative effects of text-image semantic cues for gaze estimation. More specifically, coarse directions (pseudo labels) can be readily represented using natural language and subsequently encoded through pre-trained CLIP encoders, thereby enhancing the discriminative embeddings with robust guidance ability.
While linguistic prompts offer effective practice for expressing high-level and coarse-grained directions, devising appropriate prompts from facial/eye images can pose a significant challenge, as articulating the nuances between input images in words can often be a difficult undertaking. Unlike image classification tasks, the generation of meaningful language descriptions (such as gaze direction prompt) using the full-fledged BLIP [43] is inherently impractical as the training of BLIP fully ignores the samples of gaze faces, thereby impeding the applicability of CLIP in the field of gaze estimation.
To address the aforementioned challenges, this paper delves into the intricate process of formulating suitable prompts and proceeds to carefully design a language-vision protocol (GazeCLIP). This protocol aims at enhancing the semantic space of the gaze estimation model. Specifically, we pre-define the text prompt as “A photo of a face gazing [class]”, where [class] represents a set of base direction(front, up, down, left, right). In order to make the constructed prompt conform to the language prior knowledge of the CLIP model, we directly use the zero-shot CLIP model to assign a direction for each image. We calculate the similarity between each input image and the sentence composed of the five directions mentioned above, and take the direction with the highest similarity as the text input. On top of the generated prompt-image pairs, we presence a cross-attention-based module to align the representations from two heterogeneous modalities. Following the fine-grained fusion of multi-modal features through transformer blocks and cross-attention condensers, the resulting refined image embedding are combined with the original image embedding together to act as the input of final regression head. As depicted in Fig. 1, Unprecedentedly, our emphasis lies on harnessing the guidance provided by textual signals, rather than solely struggling to design image encoders.
We evaluate our proposed model on three challenging public dataset benchmarks for gaze estimation. Extensive experiments demonstrate that GazeCLIP achieves state-of-the-art performance in all three datasets. Particularly, compared with existing approaches, the GazeCLIP gains significant improvement with an angular error decrease of 0.5° on average(4.1° → 3.6°, 7.7° → 7.3°, 5.3° → 4.7°). In a nutshell, the overall contributions of our GazeCLIP are threefold:
  • In Use. A novel text-guided gaze estimation method is proposed in this paper, namely GazeCLIP, to enhance the performance of gaze estimation, which is inspired by strong generalization capability of CLIP model in various downstream vision tasks. To the best of our knowledge, this is the first effort to distil rich knowledge from full-fledged pre-trained language-vision model for guiding the learning of gaze estimation network.
  • Fine-grained Learning. In this study, we introduce a cross-attention condenser designed to finely recalibrate visual and text representations. This mechanism facilitates the nuanced alignment of image features with the semantic guidance embedded in textual signals, enhancing the learning quality of gaze features.
  • SOTA Performance. Extensive experiments conducted on three prominent datasets unequivocally demonstrate the superior performance of our framework. Our innovative approach yields a noteworthy enhancement, manifesting as an average reduction of 0.5° (equivalent to a substantial 9.3% improvement) in angular error.
2 Related Work
2.1 Gaze Estimation Evolution
Early approaches attempt to achieve gaze estimation by predicting a mass of points on a 2D screen[55]. Owing to significant variations in the positions of cameras across different devices, the generalization capability of the 2D gaze estimation model is notably constrained. To mitigate this issue, a series of 3D gaze estimation approaches[11, 12, 13][17][23, 24, 25] are proposed to infer person gaze in the real world by enriching geometric information, such as varying shooting conditions. In contrast to geometric-based approaches, appearance-based models have garnered escalating attention within the research community. This is attributed to the capability of gaze regressors of this type to utilize images captured by conventional cameras as input. Benefiting from the huge success of deep learning[45][56], the performance of gaze estimation algorithms has been substantially improved [22]. These methods establish mapping functions between raw images and gazes through the powerful nonlinear fitting ability of neural network, and can achieve high accuracy even on unconstrained settings.
Specifically, the approach [22] unprecedentedly leverages high-capacity CNN to regress gazes from the facial image inputs. Zhang et al. [17] introduce a spatial attention mechanism to emphasize on the facial areas within input scenes, effectively suppressing the noises from redundant image regions. Meanwhile, Cheng et al. [23] realize the asymmetry between two eyes and further propose an asymmetric regression backbone consisting of four CNN stems. Stimulated by atrous convolution in image classification[57], Chen et al. [24] strive to enlarge receptive fields through dilated convolutions without incurring extra computational overheads. Wang et al[25] present a unified framework including both adversarial learning and a Bayesian approach, thereby enhancing the transferability and accuracy of gaze estimation models. The CA-Net [11] delves into the coarse-to-fine estimation procedure. In particular, the coarse stage hammers at deducing a fundamental gaze direction from a facial image, whereas the subsequent step further refines the predicted direction based on the provided eye images. More recently, [51] introduce an unsupervised domain adaptation approach to estimate gaze that is not restricted by source limitations which is superior in cross-domain gaze estimation methods. In [52], cameras with diverse viewpoints are utilised to capture the subject’s image. Following that, they set out a dual viewpoint gaze estimation network. Albeit achieving impressive results, all these gaze estimators ignore the promising semantics merits provided by large-scale modern language models (e.g., BERT [32] and CLIP [18]). Therefore, there is a large room for performance improvement in gaze estimation.
2.2 Language-steering Visual Models
Recently, Large Language Models (LLMs)[30, 31, 32] trigger a new trend of researches on text-oriented feature learning for a cohort of visual tasks, such as crowd counting[18], point clouds analysis [58] and image generatio [49]. Thereinto, Contrastive Language-Image Pre-training (CLIP) [18] model stands out as particularly captivating for enhancing visual algorithms. In specific, CLIP seeks to explore the interrelationships between image and language signals by assimilating insights from a vast dataset comprising 400 million image-text pairs collected from the internet.
CLIP is trained through contrastive learning to maximize the cosine similarity between text embedding and image embedding of positive sample pairs. Such a training method breaks the long-standing heavy reliance on image labels in the visual field. The pre-trained CLIP models can chieve superior performance compared to those fully supervised methods in many classic visual tasks even in few-shot or zero-shot settings. Given the powerful transferable ability of CLIP, many studies have explored how to apply CLIP to various specific downstream tasks. Several methods [20][26, 27, 28, 29]have achieved promising performance with the help of CLIP model on 3d avatar generation, age estimation, image aesthetics assessment, semantic segmentation and so on. These methods usually apply pre-trained CLIP model as backbone, and design a language-image interaction mechanism based on specific tasks after defining different prompts which usually are general language expression of the target images in advance.
3 Methodology
In this section, we describe how we perform gaze estimation with text knowledge. Firstly, we will illustrate the CLIP model which contains rich semantic knowledge of images and texts through unsupervised pre-training. Next, we will introduce the details in GazeCLIP.
3.1 Revise for CLIP
Influenced by large models in the field of natural language processing[30, 31, 32, 33], some research has begun to study how to apply the ideas of pre-training and large models to the field of computer vision, and break away from the dependence on large amounts of label data during visual task training since unsupervised learning models will have stronger transfer capability typically. CLIP[18], which collects 400 million pairs of image and text on the internet for model pre-training, has established the connection between visual concept and language concept. All of these texts are natural language descriptions of images. Specifically, during pre-training, given a batch of image-text pairs, they are fed into the image encoder and text encoder respectively, and then the cosine similarity of the positive sample pair is maximized through contrastive learning. By mapping the features of two modalities to a joint embedding space, the pre-trained CLIP model owns strong transfer ability, enabling it to be applied well to various downstream visual tasks even in zero-shot setting. For example, as image classification, a series of class labels (e.g., ‘car,’ ‘plane’) can be constructed to a prompt template rather than a single word since a complete sentence description is used in CLIP pre-training. The most common example is “a photo of [CLASS]”. The prompt is then fed into the text encoder to create class embeddings, which are utilized for computing similarity with image embeddings for classification. When applying the CLIP model to other downstream tasks, two issues need to be considered: firstly, constructing appropriate prompt that match the image, and secondly, effectively integrating the image embeddings and text embeddings. In the original CLIP model, the focus is on the similarity between the image and the text which is still a classification task essentially. Therefore, it needs to be adjusted for different tasks. In this work, we are the first to explore how to transfer the rich image and text knowledge learned by CLIP to the task of gaze estimation.
Figure 2: Architecture of the GazeCLIP. We only use facial images as input and take image encoder and text encoder in CLIP as backbone to extract feature. During training, we fine-tune the image encoder while the parameters of the text encoder are frozen.
3.2 The GazeCLIP framework
In order to apply the CLIP model to the gaze estimation, we design GazeCLIP as shown in Fig. 2. Firstly, we use CLIP pre-trained image encoder and text encoder to obtain their respective embeddings. Next, in order to effectively fuse the features of the two modalities instead of calculating similarity as in the CLIP model, we designed a cross attention mechanism to extract refined image embeddings that respond to text, and added them to the original image embeddings to enhance image representation. Finally, the enhanced image representation will be used for regression prediction to get gaze.
Image Encoder. We choose the pre-trained ViT-B/32[34] of CLIP for image encoder. Given a face image $I^{\prime}\in\mathbb{R}^{H\times W\times C}$, the channel C is 3 as usual, $I^{\prime}$ is divided into patches with a size of 32×32.All patches will be flatten and passing through a linear layer. After adding positional embedding, the features are fed into transformer encoder layers and we can obtain a image embedding $I$ with a dimension of 512, which can be expressed as:
$I=ImageEncoder(I^{\prime}).$ (1)
Text Encoder. Due to the fact that the label in the gaze estimation is a 2-dimensional array corresponding to the pitch and yaw angles in the 3D gaze estimation, it is not appropriate to construct the prompt. We predefine the text prompt as “A photo of a face gazing [class]” to produce a universal phrase for images in gaze estimation dataset, and [class] denotes a group of primary directions (front, up, down, left, right). To guarantee that the prompt aligns with the language prior knowledge of the CLIP model, we employ the zero-shot CLIP model to allocate a direction to each image and the prompt composed of this direction has the highest cosine similarity to the current image. Meanwhile, we freeze the text encoder provided by CLIP to obtain well learned language prior knowledge. By tokenize and text encoder, here can obtain a text embedding T which is the same dimension with image embedding. This step can be defined as:
$T=TextEncoder(tokenize(prompt)).$ (2)
Cross Attention. To achieve more refined image representation by aligning image features with text features, a cross-attention mechanism is proposed to model the association between text embedding (Q) and image embedding (K, V). Before calculating the attention score, they all pass through a layer of linear projection. And the attention scoring function here we used is scaled dot-product attention. The above can be expressed as:
$\operatorname{score}(Q,K)=\frac{Q^{T}K}{\sqrt{D_{k}}}.$ (3)
$\tilde{I}\ =Matmul(softmax(score(Q,K)),V).$ (4)
Based on the new image embedding corresponding to text, the original image embedding $I$ is added with $\tilde{I}$ to produce the enhanced embedding $\bar{I}$ used for downstream regression head:
$\bar{I}\ =I+\tilde{I}.$ (5)
$gaze=MLP(\bar{I}\ ).$ (6)
Here the MLP contains three linear layers and two ReLu layers.
Optimization. Most appearance-based gaze estimation models predict 3D gaze as the gaze direction angles (yaw, pitch) in spherical coordinates. They are all consecutive values. Therefore, L1-loss and L2-loss can be used here to optimize the model for different datasets:
$\ell_{1}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-P_{i}\right|,$ (7)
$\ell_{2}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-P_{i}\right)^{2},$ (8)
Where $y_{i}$ is the ground truth, $p_{i}$ is the predict value. They are all 2D vectors.
4 Experiments
4.1 Datasets
To validate the performance of our proposed model, we conducted experiments on three well-known and challenging gaze estimation datasets, which are MPIIFaceGaze[16], EyeDiap [14], and RT-Gene[12]. All datasets are normalized as in [35] for a fair comparison. As in previous studies, we take one decimal place for the results here.
MPIIFaceGaze. MPIIGaze [22] is the most commonly considered benchmark dataset for unconstrained 3D gaze estimation. And the MPIIFaceGaze is an extension of the MPIIGaze dataset, which provides 45K face images instead of eyes from 15 subjects captured by laptop camera during their daily life over several months. As a result, it contains images with diverse backgrounds, captured at varying times and under different lighting conditions. Leave-one-subject-out cross-validation is followed as the previous works do [11][17][23][25] which means leave one subject’s data as the test set at a time.
EyeDiap. The original EyeDiap dataset is a series of video clips of 16 subjects. Here we pre-processed the data according to [36], extracting images every 15 frames from the video clips. And we only use the videos collected in screen target setting while we can obtain a face image facing the screen in this case. Due to videos of this type are not included in No.12 subject and No.13 subject, the processed EyeDiap dataset includes 16K images of 14 subjects. Since the EyeDiap dataset does not provide a standard evaluation subset, we also perform a Leave-one-out strategy on it to get robust results.
RT-Gene. The RT-Gene dataset contains 122531 images of 15 subjects with wearable eye-tracking glasses which used for get gaze label. With the help of a series of wearable devices, the RT-gene dataset including a farther distance between the camera and the person, a greater variation in head poses and gazes compared to previous in-the-wild datasets. Therefore, this is a more challenging dataset. We follow the 3-fold evaluation protocol provided by RT-Gene dataset and divide the 15 subjects into three groups.
4.2 Evaluation Metric
For 3D gaze estimation, the most widely used evaluation metric is angular gaze error where a lower error indicates a better model. Given the ground truth ${g^{*}}$ , the evaluation Metric is defined as:
[…]
4.3 Implementation Details
We implement the experiments with PyTorch. All experiments are conducted on Nvidia A5000 GPU. The input images are 224×224×3 face images after normalizing. The backbone in our model is ViT-B/32 for image while Transformer[37] for text, they all generate features of the same dimension size in CLIP pre-trained model. And the parameter of text encoder is frozen during training. We train the GazeCLIP on three datasets with 128 batch size and 50 epochs. To optimize the model, we apply Adam optimizer [38] with an initial learning rate of 1e-5. And L1 Loss is applied for MPIIFaceGaze and EyeDiap, while L2 loss for RT-Gene. In order to adjust the learning rate during training process, we adopt the scheduler of MultiStepLR strategy. The learning rate will decay after the fifth epoch and 45th epoch, with a decay rate of 0.1.
Table 1: Comparison with State-of-the-art methods. GazeCLIP achieves state-of-the-art results in all datasets with significant improvement.
Methods Input MPIIFaceGaze[16] RT-Gene [12] EyeDiap[14]
FullFace[17] Face 4.8° 10.0° 6.6°
RT-Gene[12] Eyes and head pose 4.8° 8.6° 6.4°
RT-Gene(4 ensemble)[12] Eyes and head pose 4.3° 7.7° 5.9°
Dilated-Net[24] Face and eyes 4.8° 8.3° 6.2°
CA-Net [11] Face and eyes 4.1° 8.2° 5.3°
GazeCLIP(Ours) Face 3.6° 7.3° 4.7°
Improvement - $12\%$ $5\%$ $11\%$
Figure 3: Left: Performance on MPIIFaceGaze dataset. Middle: Performance on RT-Gene dataset. Right: Performance on EyeDiap dataset.
5 Result and Analysis
Figure 4: Angular error of different subjects on MPIIFaceGaze dataset.
5.1 Comparison with the State-of-the-Arts
We first conduct an experiment to compare GazeCLIP with other state-of-the-art gaze estimation methods which are FullFace(spatial weights CNN)[17], Dilated-Net(a three streams network with dilated convolution)[24], RT-Gene( a three streams network with eyes and head pose)[12] and CA-Net(a coarse to fine adaptive network) [11]. Since RT-Gene report model ensemble is effective, here we present two results about RT-Gene (single model and 4 ensemble). Currently, the best performances in three datasets are achieved by CA-Net and RT-Gene (4 ensemble). Tab. 1 shows that GazeCLIP can achieve state-of-the-art results in all three datasets. Our method has improved by 0.5°($12\%$), 0.4°($5\%$), and 0.6°($11\%$) compared to the previous best results on three datasets, respectively. The excellent performance on three datasets demonstrates that our proposed GazeCLIP is a advantageous network architecture for gaze estimation.
Figure 5: Visualization results of our method for 3D gaze estimation. The red lines are the ground truth annotations and the blue lines are predicted results.
Figure 6: Some failure cases.
5.2 Ablation Study
In order to further validate the effectiveness of the modules in the framework, we conducted three sets of ablation experiments. The following experiments are conducted on the MPIIFaceGaze[16] dataset, the most commonly used dataset in gaze estimation.
5.2.1 Ablation Study for language knowledge
We first analysis the effectiveness of language knowledge. We verify this through two strategies. Firstly, we maintain the existing network structure and change the pre-defined prompt to “ ” as text input; The second method is to remove the branch of language knowledge from the network and only use the image encoder for regression which makes the network no longer contain any text features. The results are shown in Tab. 2. If we only using the original CLIP image encoder with fine-tune like previous research architecture that only contain image features, however, reports unsatisfactory performance. Because more complex backbone often leads to more severe over-fitting problem in gaze estimation. When we keep the original network structure unchanged and replace the defined prompt with “ ”, the performance is significantly improved with 0.6°(4.4° → 3.8°). And taking a pre-defined prompt, “A photo of a face gazing [class]”, for a general language expression of the gaze estimation images as the input text achieves the best result. These results demonstrate the necessity of text features and the effectiveness of a appropriate prompt for text input in GazeCLIP.
Table 2: Ablation study for language knowledge. The pre-defined prompt is “A photo of a face gazing [class]”.
Prompt Angular error $\Delta$
Pre-defined prompt 3.6° -
“ ” 3.8° -0.2°
- 4.4° -0.8°
5.2.2 Ablation Study for fixing different encoders
We further explore the effect of freezing image encoder and text encoder. As listed in Tab. 3, if we directly use the original CLIP model with a cross-attention mechanism and a regression head(i.e., the parameters of the image encoder and text encoder will not be optimized), we get the angular error of 8.9° which is not satisfactory. While we unfreeze the text encoder, it does not bring significant change and even worse(8.9° → 9.0°). However, when we choose to fine-tune the image encoder, here is a significant improvement with a drop form 8.9° to 3.8°. And if we try to fine-tune the text encoder as well, the angular error increased from 3.6° to 3.8°. Whether or not the image encoder is fixed, the fixed text encoder setting always performs better. We consider that the pre-trained CLIP model contains rich language priors that can encode natural languages well, and fine-tuning will break this. In addition, due to the data for CLIP pre-training come from a wide range of internet images, some special images such as in gaze estimation dataset may only account for a small part. As reported in [18], the zero-shot CLIP model can get a relative high accuracy in most image classification datasets but a simple dataset, MNIST[39]. Since the dataset for CLIP pre-training hardly contains handwritten digital images like the images in MNIST. Therefore, we need fine-tune the image encoder to cope with some particular tasks.
Table 3: Ablation Study of freezing different encoders.
Fixed image encoder Fixed text encoder Angular error
$\surd$ $\surd$ 8.9°
$\surd$ - 9.0°
- $\surd$ 3.6°
- - 3.8°
5.2.3 Ablation Study for feature fusion method
Then we study the influence of using different feature fusion methods during the training phase. In addition to using the cross attention mechanism we mentioned earlier, we also attempt two common methods: concat and add. Specifically, after obtaining image embedding and text embedding from CLIP encoders separately, we directly concatenate or add these two features. Since they have the same dimension, feature concat will double the feature dimension thus we need to modify the input dimension of the first linear layer in regression head, and feature add will not change the feature dimension, nor will cross attention mechanism. Results are shown in Tab. 4. Whether the feature fusion method is changed to concat or add, the angular error is 4.4°. Compared to the cross attention mechanism we use, it has increased by 0.8°. The results show that the cross attention mechanism is an effective way to establish an interaction between the language knowledge and image knowledge learned from the CLIP model.
Table 4: Ablation Study for feature fusion method.
Feature fusion method Angular error $\Delta$
Cross attention 3.6° -
Concat 4.2° -0.6°
Add 4.4° -0.8°
5.3 Visual results
We display some visual images of our method in Fig. 5. These images indicate that our method can perform well in various situations, including subject gender, facial appearance, lighting conditions, whether glasses are worn , etc. We also presented some cases with poor results. As shown in Fig. 6, these images all have some special characteristics, such as partially closed eyes, dim facial lighting, blurred eye areas, etc. Due to not additional using eye images or paying more attention to the eye area, our method did not achieve satisfactory results on these images and we leave this problem as future work.
Moreover, as shown in the sixth sub-figure in Fig. 6 and Fig. 4, when take the 14th subject in the MPIIFaceGaze dataset as the test set, its experimental results show the highest angular error in 15-fold cross validation even if the appearance and lighting of the images are not very different from others. We speculate that the possible reason is that the images of the 14th subject always uses one hand to hold her face, which is almost non-existent in other subjects.
6 Conclusion
In this paper, we propose a novel network based on the vision-language pre-trained model (CLIP) to get accurate gaze estimation, named as GazeCLIP. In order to apply language knowledge to gaze estimation task, we uniformly establish corresponding language expressions for all images. And a key point is how to establish the connection between the two features. To make language knowledge effective, we designed a cross attention mechanism to enhance image representation. Experiments show the GazeCLIP can achieve state-of-the-art results in three well-known gaze estimation datasets ($12\%$, $5\%$, $11\%$improvement, respectively). We further conduct experiments to assess the advantages of different modules in our method. We hope this paper will provide a new insight for future research of gaze estimation with visual-language knowledge.