### v3

# GazeCLIP: Towards Enhancing Gaze Estimation via Text Guidance

Abstract
Over the past decade, visual gaze estimation has garnered increasing attention within the research community, owing to its wide-ranging application scenarios. While existing estimation approaches have achieved remarkable success in enhancing prediction accuracy, they primarily infer gaze from single-image signals, neglecting the potential benefits of the currently dominant text guidance. Notably, visual-language collaboration has been extensively explored across various visual tasks, such as image synthesis and manipulation, leveraging the remarkable transferability of large-scale Contrastive Language-Image Pre-training (CLIP) model. Nevertheless, existing gaze estimation approaches overlook the rich semantic cues conveyed by linguistic signals and the priors embedded in CLIP feature space, thereby yielding performance setbacks. To address this gap, we delve deeply into the text-eye collaboration protocol and introduce a novel gaze estimation framework, named GazeCLIP. Specifically, we intricately design a linguistic description generator to produce text signals with coarse directional cues. Additionally, a CLIP-based backbone that excels in characterizing text-eye pairs for gaze estimation is presented. This is followed by the implementation of a fine-grained multi-modal fusion module aimed at modeling the interrelationships between heterogeneous inputs. Extensive experiments on three challenging datasets demonstrate the superiority of the proposed GazeCLIP which achieves the state-of-the-art accuracy.
[…]
123
1 Introduction
During the past decade, visual gaze estimation has garnered ever-increasing attention thanks to its widespread application scenarios, such as saliency detection [47], virtual reality[51], human-robot interaction[19, 44], medical diagnosis[4], driver fatigue estimation[53] and other areas[37, 13]. Gaze estimation aims to accurately infer the direction of a person’s gaze from an input image, typically focusing on the human face or eyes. Despite its promising practical potential applications, this visual task faces significant challenges. For instance, the utilization of diverse camera angles inevitably leads to unpredictable variations in lighting, color, and facial features, which can severely degrade performance. To struggle with this issue, several existing studies[59, 7, 5, 46, 3, 6, 1, 48] have made lots of efforts to marry the stability and robustness to regression algorithms. An additional catalyst for the thriving of investigations on gaze estimation is the increasing availability of large-scale datasets [15, 21, 16, 61, 58] which consist of varied subjects including lighting conditions, head poses and so on.
Figure 1: (a) Existing single-modal approaches directly learn gaze-oriented representations from 2D face/eye images via CNNs-based structures, whereas (b) our proposed novel GazeCLIP delves deep into the synergistic effects of text-image features.
Inspired by the huge success of deep learning in computer vision[24, 18, 40], an increasing number of researchers are shifting their focus towards modeling intricate mappings between input signals and gaze labels[58, 59, 57]. Nonetheless, these methods typically operate on single-modal facial images as input, making them susceptible to the challenge of isolating gaze-specific features amidst unrelated regions (e.g., the nose and mouth) scattered throughout the input image as different regions are prone to confuse the model on what to prioritize in learning. A general model for remote POG estimation based on video[17] evades this problem by mining geometric characteristics of the eye. Albeit the impressive progress, conventional geometry-modelling approaches exhibit several limitations such as reliance on controlled laboratory environments and poor generalization performance to new subject. This has led to a surge in the booming of appearance-based gaze estimation algorithms, harnessing the capabilities of cutting-edge deep learning models. To enhance the estimation performance, recent learning-based protocols [9, 31, 8] excel at extracting useful features from raw facial or eye images through the design of diverse high-capacity CNN models, see Figure 1(a). Furthermore, the inevitable drastic variations introduced by diverse camera views and individual characteristics (e.g. gender, skin color, and glasses) [15, 57] can easily lead the model to fail in learning highly non-nonlinear relationships, resulting in overfitting and sub-par performance. Therefore, there is a large room for further exploration of gaze regression model with enhanced learning capabilities.
Recently, a cohort of large-scale pre-traiend linguistic-vision models [34, 12, 36, 14] have emerged as dominant forces in artificial intelligence, sparking a surge in multi-modal learning. Particularly, Contrastive Language-Image Pre-Training (CLIP) [35] elicits the foremost attention within the research community and has been widely adopted for various downstream vision tasks, such as image generation and visual question answering [38, 39, 41]. Specifically, CLIP’s backbone is comprised of a set of transformer blocks and is trained using an extensive collection of image-text pairs (400 million). The key advantage of incorporating CLIP lies in its implicit enrichment of features with the nuanced semantics embedded in natural language. For example, object detection [42], semantic segmentation [50], and image classification [62] have achieved notable success by harnessing the guiding capabilities of the prominent CLIP model.
Despite the remarkable progress made in previous approaches, the field of gaze estimation has yet to fully exploit the substantial benefits of leveraging language guidance. Indeed, the task of regressing raw face/eye images into continuous gaze vectors presents an intractable challenge. Therefore, a "divide-and-conquer" approach naturally emerges, aiming to initially identify coarse-grained gaze directions and then utilize the coarse-granularity pseudo labels to guide fine-grained gaze estimation. Moreover, considering the potential decrease in estimation accuracy when the target gaze direction deviates from the camera’s central axis compared to direct gaze towards the camera, we posit that the estimation accuracy can be notably enhanced through the guidance of coarse-grained pseudo-directions. These coarse directions, represented as natural language prompts, can be encoded through pre-trained CLIP encoders, thereby enhancing the discriminative embeddings with robust guidance ability.
While linguistic prompts effectively convey high-level and coarse-grained directions, devising appropriate prompts from facial/eye images poses a significant challenge, as articulating the nuances between input images in words is difficult. Unlike image classification tasks, the generation of meaningful language descriptions (such as gaze direction prompt) using the full-fledged BLIP [26] is inherently impractical as the training of BLIP fully ignores gaze face samples, thereby impeding the applicability of CLIP in gaze estimation. Due to the particularity of the gaze estimation dataset, natural languages generated using BLIP will always be "A man’s/woman’s face", but what we need for our approach is a rough direction guide, and it will not appear in common language descriptions.
To address the aforementioned challenges, this paper delves into the intricate process of formulating suitable prompts and proceeds to carefully design a language-vision protocol (GazeCLIP). This protocol aims at enhancing the semantic space of the gaze estimation model. Specifically, we pre-define the text prompt as “A photo of a face gazing [class]”, where [class] represents a set of base direction(front, up, down, left, right). To make the constructed prompt conform to the language prior knowledge of the CLIP model, we directly use the zero-shot CLIP model to determine a direction for each image. We calculate the similarity between each input image and the sentence composed of the five directions mentioned above, and take the direction with the highest similarity as the text input. The main reason for this is that we hope to build a natural language that conforms to prior knowledge of the CLIP model. On top of the generated prompt-image pairs, we present a cross-attention-based module to align the representations from two heterogeneous modalities. Following the fine-grained fusion of multi-modal features through transformer blocks and cross-attention condensers, the resulting refined image embedding are combined with the original image embedding together to act as the input of final regression head. As depicted in Fig. 1, Unprecedentedly, our emphasis lies on harnessing the guidance provided by textual signals, rather than solely struggling to design image encoders.
We evaluated our proposed model on three challenging public dataset benchmarks for gaze estimation. Extensive experiments demonstrate that GazeCLIP achieves state-of-the-art performance in all three datasets. Particularly, compared with existing approaches, the GazeCLIP gains significant improvement with an angular error decrease of 0.4° on average(4.1° → 3.6°, 7.4° → 7.3°, 5.3° → 4.7°). In a nutshell, the contributions of our GazeCLIP are threefold:
  • In Use. A novel text-guided gaze estimation method is proposed in this paper, namely GazeCLIP, to enhance the performance of gaze estimation, which is inspired by strong generalization capability of CLIP model in various downstream vision tasks. To the best of our knowledge, this is the first attempt to distil rich knowledge from full-fledged pre-trained language-vision model for guiding the learning of gaze estimation network.
  • Fine-grained Learning. In this study, we introduce a cross-attention condenser designed to finely recalibrate visual and text representations. This mechanism facilitates the nuanced alignment of image features with the semantic guidance embedded in textual signals, enhancing the learning quality of gaze features.
  • SOTA Performance. Extensive experiments conducted on three prominent datasets unequivocally demonstrate the superior performance of our framework. Our innovative approach yields a noteworthy enhancement, manifesting as an average reduction of 0.4° (equivalent to a substantial 8% improvement) in angular error.
2 Related Work
2.1 Gaze Estimation Evolution
Early approaches attempt to achieve gaze estimation by predicting a mass of points on a 2D screen[23]. Due to significant variations in the positions of cameras across different devices, the generalization capability of the 2D gaze estimation model is notably constrained. To mitigate this issue, a series of 3D gaze estimation approaches[8, 15, 21, 59, 7, 5, 46] are proposed to infer person gaze in the real world by enriching geometric information, such as varying shooting conditions. In contrast to geometric-based approaches, appearance-based models have garnered escalating attention within the research community. This is attributed to the capability of gaze regressors of this type to utilize images captured by conventional cameras as input. Taking advantage of the great success of deep learning[18, 43], the performance of gaze estimation algorithms has improved substantially.
Specifically, the approach [57] unprecedentedly leverages high-capacity CNN to regress gazes from the facial image inputs. Zhang et al. [59] introduce a spatial attention mechanism to emphasize facial areas within input scenes, effectively suppressing noise from redundant image regions. Meanwhile, Cheng et al. [7] realize the asymmetry between two eyes and further propose an asymmetric regression backbone consisting of four CNN stems. Stimulated by atrous convolution in image classification[54], Chen et al. [5] strives to enlarge receptive fields through dilated convolutions without incurring extra computational overhead. Wang et al[46] present a unified framework including both adversarial learning and a Bayesian approach, thereby enhancing the transferability and accuracy of gaze estimation models. The CA-Net [8] delves into the coarse-to-fine estimation procedure. In particular, the coarse stage hammers at deducing a fundamental gaze direction from a facial image, whereas the subsequent step further refines the predicted direction based on the provided eye images. [2] began to introduce an attention mechanism to extract key features of eye images. More recently, [3] introduced an unsupervised domain adaptation approach to estimate gaze that is not restricted by source limitations, which is superior in cross-domain gaze estimation methods. In [6], cameras with various viewpoints are used to capture the subject’s image. Following that, they set out a dual viewpoint gaze estimation network. [52, 32, 11, 29, 30] are also concerned about the performance under cross-dataset settings and aim to build more generalized models for gaze estimation task. [49] has demonstrated that higher resolution images will have a positive impact on the results, even with a simple backbone network. Albeit achieving impressive results, all these gaze estimators ignore the promising semantics merits provided by large-scale modern language models (e.g., BERT [12] and CLIP [35]). Therefore, there is a large room for performance improvement in gaze estimation.
2.2 Language-steering Visual Models
Recently, Large Language Models (LLMs)[33, 34, 12] trigger a new trend of researches on text-oriented feature learning for a cohort of visual tasks, such as crowd counting[28], point clouds analysis [56] and image generatio [39]. Thereinto, Contrastive Language-Image Pre-training (CLIP) [35] model stands out as particularly captivating for enhancing visual algorithms. In specific, CLIP seeks to explore the interrelationships between image and language signals by assimilating insights from a vast dataset comprising 400 million image-text pairs collected from the internet.
CLIP is trained through contrastive learning to maximize the cosine similarity between text embedding and image embedding of positive sample pairs. Such a training method breaks the long-standing heavy reliance on image labels in the visual field. The pre-trained CLIP models can chieve superior performance compared to those fully supervised methods in many classic visual tasks even in few-shot or zero-shot settings. Given the powerful transferable ability of CLIP, many studies have explored how to apply CLIP to various specific downstream tasks. Several methods [50, 20, 27, 28, 55]have achieved promising performance with the help of the CLIP model on 3d avatar generation, age estimation, image aesthetics assessment, semantic segmentation, and so on. These methods usually apply pre-trained CLIP model as backbone, and design a language-image interaction mechanism based on specific tasks after defining different prompts which usually are general language expression of the target images in advance.
3 Methodology
In this section, we outline our approach for performing gaze estimation with text knowledge. We begin by illustrating the CLIP model, which contains rich semantic knowledge of images and texts through unsupervised pre-training. Next, we delve into the details in GazeCLIP.
3.1 Revise for CLIP
Influenced by the success of large models in the field of natural language processing[33, 34, 12, 36], research has begun to study how to apply the ideas of pre-training and large models to the field of computer vision. This shift aims to reduce dependence on large amounts of labeled data during visual task training, as unsupervised learning models typically exhibit stronger transfer capabilities. CLIP[35], exemplifies this approach, leveraging 400 million pairs of image-text data collected from the internet for pre-training. These texts serve as natural language descriptions of images. During pre-training, a batch of image-text pairs is fed into separate image and text encoders, and the contrastive learning framework maximizes the cosine similarity of positive sample pairs. By mapping features of two modalities to a joint embedding space, the pre-trained CLIP model possesses strong transferability, making it applicable to various downstream visual tasks, even in zero-shot settings. For instance, in image classification, a series of class labels (e.g., ‘car’, ‘plane’) can be constructed to a prompt template, such as “a photo of [CLASS]”, as utilized in CLIP pre-training. This prompt is then fed into the text encoder to create class embeddings, which are used to compute similarity with image embeddings for classification. When applying the CLIP model to other downstream tasks, two considerations arise: constructing appropriate prompts matching the image and effectively integrating image embeddings and text embeddings. While the original CLIP model focuses on image-text similarity, often for classification tasks, adjustments are necessary for different tasks. In this work, we are the first to explore transferring the rich image and text knowledge learned by CLIP to the task of gaze estimation.
Figure 2: Architecture of the GazeCLIP. We only use facial images as input and take the image encoder and text encoder in CLIP as the backbone to extract feature. During training, we fine-tune the image encoder while the parameters of the text encoder are frozen.
3.2 The GazeCLIP framework
In order to apply the CLIP model to the gaze estimation, we propose GazeCLIP, as shown in Fig. 2. Initially, we use the CLIP pre-trained image encoder and text encoder to obtain their respective embeddings. Subsequently, instead of simply computing similarity as in the CLIP model, we employ a cross-attention mechanism to effectively fuse the features of the two modalities. This mechanism extracts refined image embeddings that respond to text and combines them with the original image embeddings to enhance the image representation. Finally, enhanced image representation is used for regression prediction to estimate gaze directions.
Image Encoder. We choose the pre-trained ViT-B/32[14] of CLIP for image encoder. Given a face image $I^{\prime}\in\mathbb{R}^{H\times W\times C}$, the channel C is 3 as usual, $I^{\prime}$ is divided into patches with a size of 32×32.All patches will be flatten and passing through a linear layer. After adding positional embedding, the features are fed into transformer encoder layers and we can obtain a image embedding $I$ with a dimension of 512, which can be expressed as:
$I=ImageEncoder(I^{\prime}).$ (1)
Text Encoder. Due to the fact that the label in the gaze estimation is a 2-dimensional array corresponding to the pitch and yaw angles in the 3D gaze estimation, it is not appropriate to construct the prompt. We pre-define the text prompt as ’A photo of a face gazing [class]’ to produce a universal phrase for images in the gaze estimation dataset, and [class] denotes a group of primary directions (front, up, down, left, right). To guarantee that the prompt aligns with the language prior knowledge of the CLIP model, we employ the zero-shot CLIP model to allocate a direction to each image and the prompt composed of this direction has the highest cosine similarity to the current image. Meanwhile, we freeze the text encoder provided by CLIP to obtain well-learned language prior knowledge. By tokenization and text encoder, here we can obtain a text embedding T which is the same dimension as the image embedding. This step can be defined as:
$T=TextEncoder(tokenize(prompt)).$ (2)
Cross Attention. To achieve more refined image representation by aligning image features with text features, a cross-attention mechanism is proposed to model the association between text embedding (Q) and image embedding (K, V). Before calculating the attention score, they all pass through a layer of linear projection. And the attention scoring function here we used is scaled dot-product attention. The above can be expressed as:
$\operatorname{score}(Q,K)=\frac{Q^{T}K}{\sqrt{D_{k}}}.$ (3)
$\tilde{I}\ =Matmul(softmax(score(Q,K)),V).$ (4)
Based on the new image embedding corresponding to text, the original image embedding $I$ is added with $\tilde{I}$ to produce the enhanced embedding $\bar{I}$ used for downstream regression head:
$\bar{I}\ =I+\tilde{I}.$ (5)
$gaze=MLP(\bar{I}\ ).$ (6)
Here the MLP contains three linear layers and two ReLu layers.
Optimization. Most appearance-based gaze estimation models predict 3D gaze as the gaze direction angles (yaw, pitch) in spherical coordinates. They are all consecutive values. Therefore, L1-loss and L2-loss can be used here to optimize the model for different datasets:
$\ell_{1}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-P_{i}\right|,$ (7)
$\ell_{2}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-P_{i}\right)^{2},$ (8)
Where $y_{i}$ is the ground truth, $p_{i}$ is the predict value. They are all 2D vectors.
4 Experiments
4.1 Datasets
To validate the performance of our proposed model, we conducted experiments on three well-known and challenging gaze estimation datasets, which are MPIIFaceGaze[58], EyeDiap [16], and RT-Gene[15]. All datasets are normalized as in [60] for a fair comparison. As in previous studies, we take one decimal place for the results here.
MPIIFaceGaze. MPIIGaze [57] is the most commonly considered benchmark dataset for unconstrained 3D gaze estimation. And the MPIIFaceGaze is an extension of the MPIIGaze dataset, which provides 45K face images instead of eyes from 15 subjects captured by laptop camera during their daily life over several months. As a result, it contains images with diverse backgrounds, captured at varying times and under different lighting conditions. Leave-one-subject-out cross-validation is followed as the previous works do [8, 59, 7, 46] which means leaving one subject’s data as the test set at a time.
EyeDiap. The original EyeDiap dataset is a series of video clips of 16 subjects. Here we pre-processed the data according to [10], extracting images every 15 frames from the video clips. And we only use the videos collected in screen target setting while we can obtain a face image facing the screen in this case. Due to videos of this type are not included in No.12 subject and No.13 subject, the processed EyeDiap dataset includes 16K images of 14 subjects. Since the EyeDiap dataset does not provide a standard evaluation subset, we also perform a Leave-one-out strategy on it to get robust results.
RT-Gene. The RT-Gene dataset contains 122531 images of 15 subjects. With the help of a series of wearable devices, the RT-gene dataset including a farther distance between the camera and the person, a greater variation in head poses and gazes compared to previous in-the-wild datasets. Therefore, this is a more challenging dataset. We follow the 3-fold evaluation protocol provided by RT-Gene dataset and divide the 15 subjects into three groups.
4.2 Evaluation Metric
For 3D gaze estimation, the most widely used evaluation metric is angular gaze error where a lower error indicates a better model. Given the ground truth ${g^{*}}$ , the evaluation metric is defined as:
[…]
4.3 Implementation Details
We implement the experiments with PyTorch. All experiments are conducted on Nvidia A5000 GPU. The input images are 224×224×3 face images after normalizing. The backbone in our model is ViT-B/32 for image while Transformer[45] for text, they all generate features of the same dimension size in CLIP pre-trained model. And the parameter of text encoder is frozen during training. We train GazeCLIP on three datasets with 128 batch sizes and 50 epochs. To optimize the model, we apply Adam optimizer [22] with an initial learning rate of 1e-5. And L1-Loss is applied for MPIIFaceGaze and EyeDiap, while L2-loss for RT-Gene. In order to adjust the learning rate during training process, we adopt the scheduler of MultiStepLR strategy. The learning rate will decay after the fifth epoch and 45th epoch, with a decay rate of 0.1.
Table 1: Comparison with State-of-the-art methods. GazeCLIP achieves state-of-the-art results in all datasets with significant improvement.
Methods Input MPIIFaceGaze[58] RT-Gene [15] EyeDiap[16]
FullFace[59] Face 4.8° 10.0° 6.6°
RT-Gene[15] Eyes and head pose 4.8° 8.6° 6.4°
RT-Gene(4 ensemble)[15] Eyes and head pose 4.3° 7.7° 5.9°
Dilated-Net[5] Face and eyes 4.8° 8.3° 6.2°
Gaze360[21] Face and eyes 4.1° - 5.3°
CA-Net [8] Face and eyes 4.1° 8.2° 5.3°
AGE-Net [2] Face and eyes 4.1° 7.4° -
GazeCLIP(Ours) Face 3.6° 7.3° 4.7°
Improvement - $12\%$ $1\%$ $11\%$
5 Result and Analysis
Figure 3: Angular error of different subjects on MPIIFaceGaze dataset.
5.1 Comparison with the State-of-the-Arts
We first conduct an experiment to compare GazeCLIP with other state-of-the-art gaze estimation methods. Since RT-Gene report model ensemble is effective, here we present two results about RT-Gene (single model and 4 ensemble). Currently, the best performance in three datasets is achieved by CA-Net and AGE-Net. Tab. 1 shows that GazeCLIP can achieve state-of-the-art results in all three datasets. Our method has improved by 0.5°($12\%$), 0.1°($1\%$), and 0.6°($11\%$) compared to the previous best results on three datasets, respectively. The excellent performance on three datasets demonstrates that our proposed GazeCLIP is an advantageous network architecture for gaze estimation.
Figure 4: Visualization results of our method for 3D gaze estimation. The red lines are the ground truth annotations and the blue lines are predicted results.
Figure 5: Some failure cases.
5.2 Ablation Study
In order to further validate the effectiveness of the modules in our framework, we conduct three sets of ablation experiments. The following experiments are carried out on the MPIIFaceGaze[58] dataset, the most widely used dataset in gaze estimation.
5.2.1 Ablation Study for language knowledge
We first analyze the effectiveness of language knowledge through two strategies. Firstly, we maintain the existing network structure and change the pre-defined prompt to empty strings as text input; Secondly, we remove the branch of language knowledge from the network and only use the image encoder for regression, eliminating any text features from the network. This allows us to verify whether relying solely on pre-trained image encoders can achieve satisfactory performance. The results are summarized in Tab. 2. However, if we only keep the original CLIP image encoder as in previous research architectures that contain only visual features, reports unsatisfactory performance. Since a more complex backbone often leads to more severe over-fitting problem in gaze estimation when the dataset size is relatively small to the pre-trained dataset in CLIP. Interestingly, when we keep the original network structure unchanged and replace the defined prompt with empty strings, the performance is significantly improved by 0.6 ° (4.4 ° 3.8 °). And taking a pre-defined prompt, “A photo of a face gazing [class]”, for a general language expression of the gaze estimation images with rough directions as the input text achieves the best result. These results demonstrate the significance for text features and the effectiveness of an appropriate design of prompts in our framework.
Table 2: Ablation study for language knowledge. The pre-defined prompt is “A photo of a face gazing [class]”.
Prompt Angular error $\Delta$
Pre-defined prompt 3.6° -
empty strings 3.8° -0.2°
No text input 4.4° -0.8°
5.2.2 Ablation Study for fixing different encoders
We further explore the effect of freezing image encoder and text encoder. As listed in Tab. 3, if we directly use the original CLIP model with a cross-attention mechanism and a regression head(i.e., the parameters of the image encoder and text encoder will not be optimized), we get the angular error of 8.9°. While we unfreeze the text encoder, it does not bring significant change and is even worse(8.9° → 9.0°). However, when we choose to fine-tune the image encoder, here is a significant improvement with a drop from 8.9° to 3.8°. And if we try to fine-tune the text encoder as well, the angular error increases from 3.6 ° to 3.8 °. Whether or not the image encoder is fixed, the fixed text encoder setting always performs better. We consider that the pre-trained CLIP model contains rich language priors that can encode natural languages well, and fine-tuning will break this. Furthermore, since the data for CLIP pre-training is drawn from a diverse range of internet images, certain characteristics present in gaze estimation datasets may be underrepresented. As reported in [35], the zero-shot CLIP model can get a relative high accuracy in most image classification datasets but a simple dataset, MNIST[25], since the dataset for CLIP pre-training hardly contains handwritten digital images like the images in MNIST. Therefore, we need fine-tune the image encoder to cope with some particular tasks.
Table 3: Ablation Study of freezing different encoders.
Fixed image encoder Fixed text encoder Angular error
$\surd$ $\surd$ 8.9°
$\surd$ - 9.0°
- $\surd$ 3.6°
- - 3.8°
5.2.3 Ablation Study for feature fusion method
Next, we study the influence of using different feature fusion methods during the training phase. In addition to using the cross-attention mechanism mentioned earlier, we also explore two common methods: concat and add. Specifically, after obtaining image embedding and text embedding from CLIP encoders separately, we directly concatenate or add these two features. Since they have the same dimension, feature concat will double the feature dimension. Then we need to modify the input dimension of the first linear layer in regression head, and feature add will not change the feature dimension, nor will cross-attention mechanism. The results are shown in Tab. 4. Whether the feature fusion method is changed to concat or add, the angular error is 4.4°. Compared to the cross-attention mechanism we use, it has increased by 0.8°. The findings demonstrate that the cross-attention mechanism serves as an effective method to establish an interaction between language knowledge and image knowledge acquired from the CLIP model.
Table 4: Ablation Study for feature fusion method.
Feature fusion method Angular error $\Delta$
Cross attention 3.6° -
Concat 4.2° -0.6°
Add 4.4° -0.8°
5.3 Visual results
We display some visual images of our method in Fig. 4. These images indicate that our method can perform well in various situations, including subject gender, facial appearance, lighting conditions, whether glasses are worn , etc. We also presented some cases with poor results. As shown in Fig. 5, these images all have some special characteristics, such as partially closed eyes, dim facial lighting, blurred eye areas, etc. We attribute the unsatisfactory performance in these instances to our method’s lack of additional utilization of eye images or focused attention on the eye area. Addressing this limitation remains an area for future investigation.
Moreover, as shown in the sixth subfigure in Fig. 5 and Fig. 3, when taking the 14th subject in the MPIIFaceGaze dataset as the test set, its experimental results show the highest angular error in 15-fold cross-validation even if the appearance and lighting of the images are not very different from others. We posit that a potential explanation lies in the consistent depiction of the 14th subject holding her face with one hand, an occurrence largely absent in images of other subjects.
6 Conclusion
In this paper, we propose a novel network, GazeCLIP, based on the vision-language pre-trained model (CLIP) to get accurate gaze estimation. We leverage language knowledge by establishing uniform language expressions for all images, aiming to enhance the effectiveness of gaze estimation. A crucial aspect of our approach is establishing the connection between visual and textual features. To achieve this, we design a cross-attention mechanism to enhance image representation. Experimental results demonstrate that GazeCLIP achieves state-of-the-art performance across three well-known gaze estimation datasets, with improvements of ($12\%$, $1\%$, $11\%$ respectively).Furthermore, we conduct experiments to assess the advantages of different modules in our method. We believe this paper provides valuable insights for future research in gaze estimation leveraging visual-language knowledge.